OpenStudy (anonymous):

Explain a method of determining the correct degree and classification of a polynomial. Why is this polynomial (4x^2y + 5xy) classified as a 3rd degree binomial? and also i need to know how to determine a perfect square trinomial.

OpenStudy (anonymous):

this will help me review for my discussion with my teacher

OpenStudy (anonymous):

@Brainybeauty

OpenStudy (anonymous):

This is a third degree binomial. It is third degree because it has three variables in its biggest term (x, x and y in 4x^2y) It is a binomial because it has two terms.

OpenStudy (anonymous):

If there are one term, it is a monomial. If there are two terms, it's a binomial. If there are three terms, it's a trinomial. And if there are many terms, it is a polynomial.

OpenStudy (anonymous):

And if you want to find out a square term, square x+1 \[(x+1)^{2} = (x+1)(x+1) = x ^{2} + x + x + 1\]

OpenStudy (anonymous):

\[= x ^{2}+2x +1\]

OpenStudy (anonymous):

You can make other perfect squares like (x+2) = x^2 +4x + 4 And of x+3, x+4, x+5 etc

OpenStudy (anonymous):

So the they are perfect squares as well as trinomials because they have three terms ( Eg. x^2 + 6x + 9)

OpenStudy (anonymous):

Anything else? @989Ƒŕĕǻҝ